OpenStudy (anonymous):

at the party 2/5 of the pizzas ordered had pepperoni .The kids ate only 1/3 of the pepperoni pizzas, while the parents ate all of the remaining pizza. How much of all the pizza ordered was eaten by the kids ?

OpenStudy (turingtest):

'of' means 'times' in math, so\[\frac2 5\times\frac1 3\]

OpenStudy (anonymous):

options are a.11/15 b.1/2 c.3/8 d.2/15 ?

OpenStudy (turingtest):

just multiply straight across with fractions\[\frac a b\times\frac c d=\frac{ac}{bd}\]and the answer should be clear

OpenStudy (anonymous):

c

OpenStudy (anonymous):

a?